OpenStudy (anonymous):

how would you solve: (square root 3x+4)-( square root 2x-4)=2

OpenStudy (anonymous):

start with \[\sqrt{3x+4}=2+\sqrt{2x-4}\] and then square both sides

OpenStudy (anonymous):

that will give you \[3x+4=4+(2x-4)+4\sqrt{2x-4}\] now put everything without a radical on the left and square again

OpenStudy (anonymous):

you get \[x-4=4\sqrt{2x-4}\] before you square

OpenStudy (anonymous):

if you square the right you would have to foil right?

OpenStudy (anonymous):

lol yeah, foil i.e. multiply

OpenStudy (anonymous):

lol at the image

OpenStudy (anonymous):

ok cool ty :) that helped alot

OpenStudy (anonymous):

don't forget to check for wrong answers

OpenStudy (anonymous):

yw